The measure of an angle is 15 less than 2 times the other angle. If the two angles are complementary angles, find the measure of each angle.
Answer by mananth(16946)

if one angle is x
the other angle = 2x-15
..
They are complementory angles
x+2x-15=90
3x-15 =90
3x=90+15
3x=105
x=105/3
x=35
the other angle = 55
...
m.ananth@hotmail.ca